OpenStudy (anonymous):

if you have a quadratic graph, how do you find the domain and range and the intercepts? what are the equations?

OpenStudy (anonymous):

I don't know much about quadratic graphs, but this night help you: http://www.purplemath.com/modules/grphquad.htm

OpenStudy (anonymous):

i can i tell whether the graph is a function or not?

OpenStudy (anonymous):

*how can i...

OpenStudy (anonymous):

Quadratic graphs are in the form \[ax^2 + bx + c\] The domain is all Real Numbers. Now for the range.. If a > 0 then the range will be from the vertex to positive infinity or \[\large range = [f(\frac{-b}{2a}), \infty)\] If a < 0 then the range is from negative infinity to the vertex or \[\large range = (-\infty, f(\frac{-b}{2a})]\]

OpenStudy (anonymous):

nevermind i had it right.

OpenStudy (anonymous):

omg thank you!!!

OpenStudy (anonymous):

if you want the x intercepts.. set the quadratic equation equal to 0.. If you want the y-intercepts plug in x to be 0.